OpenStudy (anonymous):

|f(x)|=4-(x+1)^2 Please help

OpenStudy (helder_edwin):

what do have to with this??

OpenStudy (anonymous):

graph it

OpenStudy (helder_edwin):

do u have to graph \(f(x)=4-(x+1)^2\), or given this graph \(y=|f(x)|\)??

OpenStudy (anonymous):

yea the second one i think it just says abs value f(x) with that as f(x)

OpenStudy (helder_edwin):

OK. f(x) is a parabola \[ \large y=4-(x+1)^2 \] \[ \large y-4=-(x+1)^2 \] with vertex at (-1,4) that opens downwards (due to the sign).

OpenStudy (helder_edwin):

 something like this

OpenStudy (anonymous):

wait it goes below the y axis?

OpenStudy (helder_edwin):

now finally \(y=|f(x)|\). the absolute value "folds" the plane at the x-axis. and everything ends up laying on the upper half of the plane

OpenStudy (helder_edwin):